In a sample of 80 new residents who received a coupon for a two month free membership to a country club, there were 30 people who actually took advantage of the free two month membership. When testing the hypothesis that at least 30% of the people that receive the coupon actually use it, the test statistic is 1.46. What is the p-value? (please round your answer to 4 decimal places)
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.30

Ha : p >0.30

Test statistic = z =1.46

P(z >1.46 ) = 1 - P(z < 1.46) = 1-0.9279=0.0721 using z table

P-value = 0.0721